The width of a rectangle is one third of the legnth. If the perimeter is 96, Find the width.
Answer by arunpaul(104)

Let length = L
Breadh = B = L/3
as per question 2(L+B)=96
2(L+L/3)=96
or 2/3(3L+L)=96 unit
or 8L=96*3
or L = 36 unit
and B= 36/3=12 unit
(Ans)